OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation? x^4 – 41x^2 = –400

OpenStudy (anonymous):

I need to show the work, I already know the answer is 4, -4, 5, -5

OpenStudy (anonymous):

Substitute \(u=x^2\) to yield the quadratic \(u^2-41u+400=u^2-16u-25u+400=u(u-16)-25(u-16)=(u-25)(u-16)\). If this product equals \(0\), then \(u\in\{16,25\}\). If \(u=16\), then we back-substitute \(u=x^2\) yielding \(x^2=16\), i.e. \(x=\pm4\). Similarly, \(x^2=25\) and thus \(x=\pm5\).

OpenStudy (anonymous):

Hmm.. Most of that doesn't look familiar to me... this is Algebra 2

OpenStudy (anonymous):

Or wait, I guess I am just confused about the "u" thing. Why substitute that in and not just use x?

OpenStudy (anonymous):

If we let \(u=x^2\), we can rewrite the equation as \(u^2 - 41u = -400\), which is then equivalent to \(u^2-41u+400=0\). If you're in Algebra II you should understand how to factor this... it yields \((u-16)(u-25)=0\). If the product of these two factors is \(0\), then it follows logically that either one of them must be \(0\) -- so: \[u-16=0\\u=16\\\lor\\u-25=0\\u=25\]By substituting back \(u=x^2\), we now have:\[x^2=16\\x=\pm\sqrt{16}=\pm4\\\lor\\x^2=25\\x=\pm\sqrt{25}=\pm5\]

OpenStudy (anonymous):

We substitute so that it's more readily visible as a quadratic of u, which we know how to solve by factoring.

OpenStudy (anonymous):

OHHHhhhh I see, A million thanks for your help! :)

OpenStudy (anonymous):

oh and is "u" like commonly used in this type of case or was it just your choice of letter?